3 ( log 5 - log 3 ) - ( log 5 - 2 log 6 ) = 2 - log n , find n.

Answer by nerdybill(7384)

3 ( log 5 - log 3 ) - ( log 5 - 2 log 6 ) = 2 - log n , find n.
.
3 ( log 5 - log 3 ) - ( log 5 - 2 log 6 ) = 2 - log n
3log 5 - 3log 3 - ( log 5 - 2 log 6 ) = 2 - log n
log 5^3 - log 3^3 - ( log 5 - 2 log 6 ) = 2 - log n
log 125 - log 27 - ( log 5 - 2 log 6 ) = 2 - log n
log 125 - log 27 - ( log 5 - log 6^2 ) = 2 - log n
log 125 - log 27 - ( log 5 - log 36 ) = 2 - log n
log 125 - log 27 - log 5 + log 36 = 2 - log n
log 125/27 - log 5 + log 36 = 2 - log n
log 125/(27*5) + log 36 = 2 - log n
log (125*36)/(27*5) = 2 - log n
log (125*36)/(27*5) = 2 - log n
log (125*36)/(27*5) - 2 = -log n
-[log (125*36)/(27*5) - 2] = log n
10^(-[log (125*36)/(27*5) - 2]) = n
10^(-[log (4500)/(27*5) - 2]) = n
10^(-[log 900/27 - 2]) = n
10^(-[log 100/3 - 2]) = n
10^(-[1.52287874528 - 2]) = n
10^(0.4771212547196624372950) = n
3 = n